# Tree-Like Shortcuttings of Trees

### v1

@@ 1 Introduction
Given a tree $T=(V,E)$ and an integer $k\geq 1$, a tree shortcutting of $T$ with hop-diameter $k$ is a graph $G=(V,E^{\prime})$ such that for every two vertices $u,v\in V$, there is a path in $G$ consisting of at most $k$ edges such that $d_{G}(u,v)=d_{T}(u,v)$, where $d_{G}(u,v)$ (resp., $d_{T}(u,v)$) represents the distance between $u$ and $v$ in $G$ (resp., $T$). Besides achieving small (ideally constant) hop-diameter, the most basic parameter of a tree shortcutting is its number of edges, $|E^{\prime}|$. The problem of constructing sparse tree shortcuttings with small hop-diameter has been studied extensively (under different names and settings), since the pioneering works of [Yao82, Cha87, AS87, BTS94]. The first tradeoff was given by Yao [Yao82], who studied the problem of computing range queries on paths. Later, Chazelle [Cha87] showed how to extend these tradeoffs from paths to arbitrary trees. Few years after, Alon and Schieber [AS87] studied the problem of computing semigroup product along paths and trees. Given a tree $T$ whose vertices are elements of a semigroup, the goal is to preprocess it into a succinct data structure so that every subsequent semigroup product query can be answered using as few as possible operations. In the terminology of tree shortcutting, they showed that for any $n$-point path, one can get a shortcutting with hop-diameter $k$ and $O(n\alpha_{k}(n))$ edges. Here, $\alpha_{k}(n)$ is a very slowly-growing inverse Ackermann function defined in Section˜2. In particular, their tradeoff implies that one can get hop-diameters $2$, 33, and $4$, with $O(\log{n})$, $O(\log\log{n})$, and $O(n\log^{*}{n})$ edges, respectively. For trees, their hop-diameter grows by a factor of 2: their tree shortcutting achieves hop-diameter $2k$ using $O(n\alpha_{k}(n))$ edges. They also showed that for a line metric the tradeoff is tight. Some of the applications of shortcuttings mentioned in [AS87] include finding the max-flow in a multiterminal network, verifying MST, and updating an MST after increasing the cost of one of its edges. The problem of problem of achieving an optimal tradeoff between hop-diameter and sparsity was settled by Bodlaender et al. [BTS94], who gave an optimal tradeoff for trees matching the lower bound of [AS87] for paths.
These constructions were successfully used in various applications and were also lifted to other graph families. Most notably perhaps, a seminal work of Arya et al. [ADM+95] shows that every $n$-point Euclidean metric with a constant dimension admits a $(1+\epsilon)$-spanner with hop-diameter $k$ and $O_{\epsilon}(n\alpha_{k}(n))$ edges. Spanners are formally defined in Section˜2. Roughly speaking, spanners generalize the notion of tree shortcutting: Given a tree $T$, a tree shortcutting of $T$ is a stretch-$1$ spanner of the tree metric induced by $T$. Filtser and Le [FL22b] used tree shortcutting to achieve low-treewidth embedding of planar graphs. (See [CFKL20, CLPP23, CCC+25] for additional results on low-treewidth embedding of planar and minor-free graphs and metrics.) In particular, the central part of the embedding result of [FL22b] is a low-treewidth shortcutting for trees. Another example is the compact routing scheme by Kahalon et al. [KLMS22]. (See Section˜2 for a definition of compact routing.) Their routing scheme operates on top a network which is a low-treewidth tree shortcutting with hop-diameter 2.
Once the shortcutting has been computed, it serves as a “proxy” overlay network, on which the computation can proceed, which gives rise to huge savings in a number of quality measures, including global and local space usage, as well as in various notions of running time, which change from one setting to another. In some applications where limiting the hop-distances of paths is crucial, such as in some routing schemes, road and railway networks, and telecommunication, we might need to minimize the hop-distances; for example, imagine a railway network, where each hop in the route amounts to switching a train – how many of us would be willing to use more than, say, 4 hops? Likewise, what if each hop amounts to traversing a traffic light, wouldn’t we prefer routes that minimize the number of traffic lights? In such cases, the designer of the system, or its users, might not be content with super-constant hop-distances, or even with a large constant, and it might be of significant value to achieve as small as possible hop-distances. Motivated by such practical considerations, we are primarily interested in values of hop-diameter $k$ that “approach” 1, mainly $k=2,3,4,$... as there is no practical need in considering larger values of $k$.
@@
The fundamental drawback of sparse tree shortcuttings is that they have dense subgraphs. In particular, all the aforementioned tree shortcutting constructions suffer from not being uniformly-sparse in the regime of constant $k$ — they have subgraphs with average degree of $\Omega(\log{n})$. This lower bound on average degree is true for hop-diameter $k=2$ and does not reduce with hop-diameter. Motivated by this, we initiate a systematic study of constant-hop tree shortcuttings that are “tree-like”. In particular we are interested in two notions that capture uniform sparsity: treewidth and arboricity. (Formal definitions are given in Section˜2.) These two graph families are prime examples of families admitting more efficient algorithms, in contrast to just sparse graphs. The key question underlying this work is the following.
Question 1.
Given an $n$-vertex tree, is there a shortcutting with constant hop diameter $k$ with arboricity/treewidth of $o(\log{n})$, and ideally close to constant? In particular, is it possible to achieve this using $k=3,4,$...?
Getting below the $\log{n}$ bound is impossible for $k=2$ due to the sparsity lower bound for shortcuttings with hop-diameter 2. A well-known construction achieves a bound of $O(\log{n})$ on sparsity, treewidth, and arboricity in this regime. Also, with $\Theta(\log{n})$ hops, one can get a constant bound on maximum degree and thus also on sparsity, treewidth, and arboricity [SE14]. We note, however, that the key focus of this paper is in the regime where the hop-diameter is constant. This is arguably the most important regime for all applications. A key insight of this work is that one can break the logarithmic bound, even for treewidth, by using as few as 3 hops. Building on this insight, we derive general nontrivial tradeoffs between the hop-diameter and the treewidth/arboricity, and demonstrate their applicability.
The aforementioned low-treewidth embedding result of Filtser and Le [FL22b] gives a tree shortcutting with hop-diameter $k=O(\log\log{n})$ and treewidth $t=O(\log\log{n})$. Using this result, they give a low-treewidth embedding of planar graphs into low-treewidth graphs. In particular, they show that if one can construct a tree shortcutting with hop-diameter $k$ and treewidth $t$, then one can embed a planar graph with diameter $D$ into a graph with treewidth $O(k\cdot t/\epsilon)$ with an additive distortion of $O(\epsilon D)$. Improving the product of $k\cdot t$ would immediately improve the bound on the treewidth of the embedding. The following question is left open in their work.
Question 2 ([Le23, FL22b]).
@@
Is there a tree shortcutting with treewidth $t$ and hop-diameter $k$ such that $k\cdot t=o((\log\log{n})^{2})$? Furthermore, is there such a tree shortcutting with a constant hop-diameter?
The compact routing scheme of Kahalon et al. [KLMS22] achieves memory bound of $O(\log^{2}{n})$ bits for routing on tree shortcuttings with hop-diameter 2. The reason why they achieve a bound of $O(\log^{2}{n})$ bits is essentially due to the fact that the shortcutting has arboricity/treewidth of $\Theta(\log{n})$. The main obstacle in improving their memory bound lies in understanding tree shortcuttings with small treewidth/arboricity. Breaking the bound of $\Theta(\log^{2}{n})$ bits is a main open question left in their work. Quoting [KLMS22]: “Whether or not one can use a spanner111tree shortcutting in our terminology of larger (sublogarithmic and preferably constant) hop-diameter for designing compact routing schemes with $o(\log^{2}n)$ bits is left here as an intriguing open question.”
Question 3 ([KLMS22]).
Given an $n$-vertex tree $T$, is there a compact routing scheme (operating on a shortcutting of $T$) with stretch $1$ which uses $o(\log^{2}{n})$ bits of space for every node and achieves an $o(\log n)$, and ideally constant hop-diameter?
@@ 1.1 Our contribution
Perhaps the main contribution of this work is a conceptual one—identifying the importance of ˜1. In Section˜1.1.1 we present new upper and lower bounds for tree shortcuttings, which answer ˜1 as well as ˜2 in the regime of hop-diameter $O(\log\log n)$. In Section˜1.1.2 we present some extensions and applications of these bounds, which in particular settle ˜3.
1.1.1 Bounds for tree shortcuttings
We provide a thorough investigation of ˜1. First, we show that one can break the $\log{n}$ barrier on treewidth already for hop-diameter 3. Section˜1.1.1 provides a general upper bound tradeoff between the hop-diameter and the treewidth, for all values of hop-diameter $k=O(\log\log n)$.
{restatable}
theoremtwUB For every $n\geq 1$ and every $k=O(\log\log{n})$, every $n$-vertex tree admits a shortcutting with hop-diameter $k$ and treewidth $O(k\log^{2/k}{n})$ for even $k$ and $O(k(\frac{\log{n}}{\log\log{n}})^{2/(k-1)})$ for odd $k\geq 3$. Remark. It is impossible to extend the result of Section˜1.1.1 to basic graph families, such as planar graphs or Euclidean metrics. See Section˜1.2 for more details.
We also prove a lower bound tradeoff between the hop-diameter and the treewidth that matches the upper bound of Section˜1.1.1 for all values of $k=O(\log\log n)$. Furthermore, we provide a lower bound for larger values of $k$, which together settle negatively ˜2, and in particular imply that the construction of [FL22b] with hop-diameter and treewidth both bounded by $O(\log\log n)$ is asymptotically optimal.
@@ 1.1.1 Bounds for tree shortcuttings
  • $\Omega(k\log^{2/k}{n})$ for even $k$ and $\Omega(k(\frac{\log{n}}{\log\log{n}})^{2/(k-1)})$ for odd $k\geq 3$, whenever $k\leq\frac{2}{\ln(2e)}\ln\log{n}$;
  • $\Omega((\log\log{n})^{2}/k)$ whenever $k>\frac{2}{\ln(2e)}\ln\log{n}$.
The construction of Section˜1.1.1 for $k=3$ has treewidth, and thus also arboricity, bounded by $O(\log{n}/\log\log{n})$. We next show that the bound on the arboricity can further be improved. In particular one can get rid of the factor $k$ in the tradeoff from Section˜1.1.1 by introducing some slack to the exponent of $\log{n}$. In particular, we prove the following theorem. {restatable}theoremarbUBtree For every two integers $n\geq 1$ and $k\geq 1$ and every $n$-vertex tree $T$, there is a shortcutting with hop-diameter $k$ and arboricity $O(\log^{12/(k+4)}{n})$. Moreover, when the height of the tree is $h$, then the arboricity is $O(h^{6/(k+4)})$.
Finally, we show an even better tradeoff between the hop-diameter and arboricity on paths. {restatable}theoremarbUBline For every $n\geq 1$ and every even $k\geq 2$, every $n$-point path admits a shortcutting with hop-diameter $k$ and arboricity $O(\alpha_{k/2+1}(n))$. In particular, Section˜1.1.1 shows that one can get arboricity $O(\log\log{n})$ with $k=4$ and $O(\log^{*}{n})$ with $k=6$. The tradeoff is asymptotically tight, due to the sparsity lower bound of [AS87], as the arboricity of any graph is at most its average degree (up to a factor of 2).
We note that all our shortcutting constructions can be implemented in time linear in their size. We will skip the implementation details as this is not the main focus of this work.
1.1.2 Extensions and applications
We next extend Section˜1.1.1, which provides a 1-spanner for line metrics with bounded hop-diameter and arboricity, to arbitrary doubling metrics, by increasing the stretch to $1+\epsilon$. {restatable}theoremarbUBdoubling Let $k$ be an even integer and let $\epsilon\in(0,1/6)$ be an arbitrary parameter. Then, for every positive integer $n$, every $n$-point metric with doubling dimension $d$ admits a $(1+\epsilon)$-spanner with hop-diameter $k$ and arboricity $\epsilon^{-O(d)}\alpha_{k/2+1}(n)$. This significantly strengthens the construction of Arya et al. [ADM+95], by providing a uniformly sparse (rather than just sparse) construction with constant hop-diameter. Specifically, for hop-diameter $k$, we transition from sparsity $\epsilon^{-O(d)}\alpha_{k}(n)$ to arboricity $\epsilon^{-O(d)}\alpha_{k/2+1}(n)$; in particular, we get arboricity $O(\log^{*}n)$ with an hop-diameter of 6.
Using the results in tree covers [CCL+23, CCL+24], we can lift this tradeoff for planar and minor-free metrics.
Theorem 1.1.
Let $k$ be an even integer and let $\epsilon\in(0,1/6)$ be an arbitrary parameter. Then, for every positive integer $n$, every $n$-point $K_{h}$-minor-free metric admits a $(1+\epsilon)$-spanner with hop-diameter $k$ and arboricity $O(\epsilon^{-3}\cdot 2^{\,h^{O(h)}/\epsilon}\cdot\alpha_{k/2+1}(n))$. Furthermore, if the metric is planar, one can construct such a spanner with arboricity $O(\epsilon^{-3}\cdot\alpha_{k/2+1}(n))$.
Using the results in [BFN19, MN07], we obtain the following theorem for spanners of general metrics.
Theorem 1.2.
Let $k$ be an even integer, $r\geq 1$ be an arbitrary parameter. Then, for every positive integer $n$, every $n$-point metric $M_{T}$ admits an $O(n^{1/r}\log^{1-1/r}{n})$-spanner with hop-diameter $k$ and arboricity $O(r\cdot\alpha_{k/2+1}(n))$. Alternatively, $M_{T}$ admits an $O(r)$-spanner with hop-diameter $k$ and arboricity $O(r\cdot n^{1/r}\cdot\alpha_{k/2+1}(n))$.
In Section˜5, we use the construction from Section˜1.1.1 to devise a constant-hop routing scheme for tree metrics with stretch 1 and $O(\log^{2}{n}/\log\log{n})$ bits per vertex. We note that our routing algorithm operates on top of a network with hop-diameter 3 and treewidth $O(\log{n}/\log\log{n})$. {restatable}theoremroutingUB For every $n$ and every $n$-vertex tree $T$, there is a 3-hop routing scheme in the fixed-port model for the metric $M_{T}$ induced by $T$ with stretch 1 that uses $O(\log^{2}{n}/\log\log{n})$ bits per vertex. This answers in the affirmative ˜3, which is the main open question from [KLMS22]. We show that the bound from Theorem˜1.2 on memory per vertex is the best one could get in tree metrics, regardless of the number of hops! In particular, we prove the following theorem. {restatable}theoremroutingLB There is an infinite family of trees $T_{n}$, $n>0$, such that any labeled fixed-port routing scheme with stretch 1 on a metric induced by $T_{n}$ has at least one vertex with total memory of $\Omega(\log^{2}{n}/\log\log{n})$ bits.
Using known tree cover constructions, we obtain a 3-hop $(1+\epsilon)$-stretch routing in doubling metrics. Given a tree cover, we construct a routing scheme on top of each of the trees in the cover. The key challenge is in obtaining a mechanism for efficiently (without increasing the memory usage) determining which tree to use for routing between any two given source and target points. We focus here on doubling metrics, where such a mechanism is already given [CCL+25], and we use it as a black-box to derive our result. However, this is a general framework that can be applied to any other metric family that admits a tree cover construction, provided that one has the mechanism to efficiently determine the right tree on top of which the routing proceeds.
@@ 1.1.2 Extensions and applications
theoremroutingUBdoubling For every $n$ and every $n$-point metric with doubling dimension $d$, there is a 3-hop routing scheme with stretch $(1+\epsilon)$ that uses $O_{\epsilon,d}(\log^{2}{n}/\log\log{n})$ bits per vertex.
This result provides the first routing scheme in Euclidean and doubling metrics, where the number of hops is $o(\log n)$, let alone as small as 3, and the labels consist of $o(\log^{2}{n})$ bits.
1.2 A natural limitation of low-treewidth shortcuttings and spanners
Section˜1.1.1 provides an upper bound tradeoff between the hop-diameter and treewidth of shortcuttings for trees. Low-treewidth shortcuttings do not exist in most other basic graph families, such as planar graphs and Euclidean and doubling metrics, and this limitation holds even regardless of any hop-diameter bound. Indeed, the $\sqrt{n}$ by $\sqrt{n}$ unweighted grid is a planar graph of treewidth $\sqrt{n}$, and any shortcutting of the grid must include the entire grid. Similarly, for the doubling metric induced by the grid, as well as for the Euclidean space induced by a 2-dimensional grid point set, any shortcutting must contain the entire grid. In fact, a similar limitation applies even when considering low-treewidth spanners of stretch larger than 1 (again even regardless of their hop-diameter). In particular, it was observed in [DFG08] that for any planar graph of treewidth $k$, any $t$-spanner must have treewidth $\Omega(k/t)$, and a similar lower bound on the spanner treewidth was extended to the families of bounded genus and apex-minor-free graphs. Building on these lower bounds, fixed parameter tractability results for these graph families were given in [DFG08] (refer also to [FGvL11] for results for bounded degree graphs). Low-treewidth Euclidean spanners were studied by Buchin et al. [BRS25], who showed that any set of $n$ points in the Euclidean plane admits a $t$-spanner of treewidth $O(n/t)$, and this tradeoff between stretch and treewidth is asymptotically optimal. Corneil et al. [CDKX25] showed that for any constants $w\geq 2$ and $c\geq 1$ there exist graphs of treewidth $w$, such that no spanning subgraph of treewidth w1w-1 can be an additive $c$-spanner of such a graph.

### v2

@@ 1 Introduction
Given a tree $T=(V,E)$ and an integer $k\geq 1$, a tree shortcutting of $T$ with hop-diameter $k$ is a graph $G=(V,E^{\prime})$ such that for every two vertices $u,v\in V$, there is a path in $G$ consisting of at most $k$ edges such that $\delta_{G}(u,v)=\delta_{T}(u,v)$, where $\delta_{G}(u,v)$ represents the distance between $u$ and $v$ in $T$. The problem of constructing sparse tree shortcuttings with small hop-diameter has been studied extensively (under different names and settings) since the pioneering works of [Yao82, Cha87, AS87, BTS94]. The optimal tradeoff is hop-diameter $k$ with $\Theta(n\alpha_{k}(n))$ edges, where $\alpha_{k}(n)$ is a very slowly-growing inverse Ackermann function, $\alpha_{2}(n)=\lceil\log{n}\rceil$, $\alpha_{3}(n)=\lceil\log\log{n}\rceil$, $\alpha_{4}(n)=\log^{*}n$, etc. (See Section˜2 for a formal definition.)
Tree shortcutting has many applications, for example, finding max-flow [Tar79], MST verification [Tar79], maintaining MST under edge weight increases [Tar79], computing semigroup product along paths and trees [Yao82, AS87], to name a few. In these applications, the shortcut graph serves as a compact data structure where $k$ determines the query time, and the number of edges is the space. Tree shortcutting is a key subroutine in constructing low-hop $(1+\epsilon)$-spanners for Euclidean [ADM+95, Sol13, NS07] and doubling metrics [CG06, KLMS22]. (See Section˜2 for a formal definition.)
@@
The fundamental drawback of sparse tree shortcuttings is that they are not uniformly-sparse — all of the aforementioned tree shortcutting constructions contain subgraphs with average degree of $\Omega(\log{n})$ for $k=2$ and $\Omega(\sqrt{n})$ for $k=3$. As $k$ increases, the (global) sparsity of the shortcutting reduces substantially, but the uniform sparsity remains $\Omega(\log{n})$. Motivated by this, we initiate a systematic study of low-hop tree shortcuttings that are “tree-like”. In particular, we are interested in two notions that capture uniform sparsity: treewidth and arboricity. (See Section˜2 for formal definitions.)
Low-treewidth shortcutting.
Low-treewidth shortcutting was introduced in the context of low-treewidth embedding of planar metrics. Specifically, Filtser and Le [FL22b] showed that if one can construct a tree shortcutting with hop-diameter $k$ and treewidth $t$, then one obtains an embedding of planar metrics into graphs of treewidth $O(kt/\epsilon)$ and additive distortion $\epsilon D$, where $D$ is the diameter of the input metric. They constructed a tree shortcutting with hop-diameter $k=O(\log\log{n})$ and treewidth $t=O(\log\log{n})$, and hence obtained an embedding with treewidth $O((\log\log)^{2}/\epsilon)$. This bound does not seem to be optimal — other works [FKS19, CCL+24a] used different embedding techniques to remove the dependency on $n$ but at the cost of a (much) higher dependency on $1/\epsilon$. If one can construct a shortcutting with hop-diameter $k=O(1)$ and treewidth $t=O(\alpha_{k}(n))$ (which matches the sparsity bound), then one gets an embedding of planar metrics with treewidth $O(\alpha_{k}(n)/\epsilon)=O(\log^{*}(n)/\epsilon)$, which is almost as good as $O(1/\epsilon)$. The linear dependency on $1/\epsilon$ is optimal [FL22b], whereas the other embedding techniques [FKS19, CCL+24a] have an inherent $\Omega(1/\epsilon^{2})$ barrier. Motivated by this observation, Filtser and Le [FL22b] asked:
Question 1 ([Le23, FL22b]).
@@
Is there a tree shortcutting with treewidth $t$ and hop-diameter $k$ such that $k\cdot t=o((\log\log{n})^{2})$? Furthermore, is it possible for $k\cdot t$ to approach $O(1)$?
In this work, we seek to answer a broader question: What is the exact tradeoff between the hop-diameter $k$ and the treewidth $t$ of tree shortcuttings? Understanding the full tradeoff would not only answer ˜1 but could also be useful for other applications that require a different hop-diameter. Specifically, in this paper, we use our construction for $k=3$ to resolve an open problem posed by Kahalon et al. [KLMS22] regarding compact routing schemes.
Low-arboricity shortcutting.
Given a set of points $P$ in an Euclidean (or doubling) space of dimension $d$, one can construct a sparse $(1+\epsilon)$-spanner with a small hop-diameter. This is achieved via tree covers. (See Section˜2 for a definition.) Known tree cover constructions for Euclidean and doubling metrics achieve stretch $1+\epsilon$ using $\epsilon^{-O(d)}$ trees [ADM+95, BFN19, CCL+24b]. By constructing a $k$-hop shortcutting $G_{T}$ for each tree $T$ in the cover $\mathcal{T}$, the union $\cup_{T\in\mathcal{T}}G_{T}$ is a $k$-hop $(1+\epsilon)$-spanner. For a constant $\epsilon$ and $d$, the sparsity is only $O(1)$ times worse than the sparsity of the shortcuttings, which is $O(\alpha_{k}(n))$. However, if $G_{T}$ has low treewidth for every tree $T$, the union $\cup_{T\in\mathcal{T}}G_{T}$ might have a very large treewidth. A recent work on low-treewidth Euclidean spanners by Buchin, Rehs, and Scheele [BRS25] showed that treewidth-$t$ spanners must have stretch $\Omega(n/t^{d/(d-1)})$. That is, if the stretch is $O(1)$, the treewidth has to be very large: $\Omega(n^{1-1/d})$.
While the treewidth grows substantially under taking union, the arboricity does not: if $G_{T}$ has arboricity at most $\beta$ for every $T\in\mathcal{T}$, then $\cup_{T\in\mathcal{T}}G_{T}$ has arboricity $|\mathcal{T}|\beta$, which is $O(\beta)$ for constant $\epsilon$ and $d$. This makes arboricity an appealing tree-like measure. Specifically, one could hope to construct a low-arboricity Euclidean (or doubling) $(1+\epsilon)$-spanner with small hop-diameter using low-arboricity tree shortcuttings.
@@ 1.1 Our contribution
Our key conceptual contribution is in identifying the power of tree-like shortcuttings of trees, in particular in the regime of very small hop-diameter. We employ the low-treewidth construction with hop-diameter 3 to improve the state-of-the-art on low-hop compact routing schemes in tree and doubling metrics. Using the low-arboricity shortcuttings, we devise a $(1+\epsilon)$-spanner with arboricity $O(\log^{*}{n})$ and hop-diameter 6 for doubling metrics. We also fully resolve ˜1 in the negative, which is our main technical contribution.
Treewidth.
Our main technical contribution is a lower bound on the treewidth $t$ of tree shortcuttings for any hop-diameter $k$, which holds even when the underlying graph is a path. This result fully resolves ˜1 in the negative.
@@
  • $t=\Omega(k\log^{2/k}{n})$ for even $k$ and $t=\Omega(k(\frac{\log{n}}{\log\log{n}})^{2/(k-1)})$ for odd $k\geq 3$, whenever $k\leq\frac{2}{\ln(2e)}\ln\log{n}$;
  • $t=\Omega((\log\log{n})^{2}/k)$ whenever $k>\frac{2}{\ln(2e)}\ln\log{n}$.
Specifically, Section˜1.1 implies that $t\cdot k=\Omega((\log\log n)^{2})$, and therefore, the artificially looking upper bound by Filtser and Le [FL22b] is indeed optimal.
We prove Section˜1.1 by constructing a large clique minor $K_{t}$ in any given shortcutting $H$ with hop-diameter $k$. Such a clique minor is a certificate that the treewidth of $G$ is at least $t$. Our lower bound instance is an $n$-vertex path, and the proof is by induction on $k$. We observe that when $k=2$, shortcutting $H$ contains the clique ${K}_{\lfloor\log{n}\rfloor+1}$ as a minor. For a higher even value of $k\geq 4$, we use the induction hypothesis for k2k-2. We divide the path into equally sized subpaths and consider two cases. If there is a subpath where every vertex has an edge going out of it, we can readily construct a large minor using this subpath. Otherwise, each subpath $P$ has at least one vertex without an edge going out of $P$. This allows us to use the induction hypothesis for k2k-2 on a graph obtained by contracting each subpath into a single vertex. The argument for odd values of $k$ is analogous. Although the method of constructing the minor is simple to describe, the analysis is much more intricate, specifically handling the dependency between $k$ and $n$ in the recurrences. See Section˜3 for details.
We also give an upper bound construction that matches the lower bound when $k=O(\log\log n)$, which arguably is the most interesting regime. Our construction is a rather natural adaptation of the known result by Filtser and Le [FL22b]. The difference is that when we shortcut a set of “cut vertices” recursively, we set the hop bound to be k2k-2. This requires some technical work; the proofs are deferred to Section˜3.1. The following theorem summarizes the construction.
@@
theoremtwUB For every $n\geq 1$ and every $k=O(\log\log{n})$, every $n$-vertex tree admits a shortcutting with hop-diameter $k$ and treewidth $O(k\log^{2/k}{n})$ for even $k$ and $O(k(\frac{\log{n}}{\log\log{n}})^{2/(k-1)})$ for odd $k\geq 3$.
In Section˜5, we use this construction to devise a 3-hop routing scheme for tree metrics with stretch 1 and $O(\log^{2}{n}/\log\log{n})$ bits per vertex. This answers the open question from [KLMS22]: “Whether or not one can use a spanner of larger (sublogarithmic and preferably constant) hop-diameter for designing compact routing schemes with $o(\log^{2}n)$ bits is left here as an intriguing open question.” In Section˜6, we prove that the bound is tight. Using tree covers, we obtain a 3-hop routing scheme with stretch $(1+\epsilon)$ that uses $\epsilon^{-\tilde{O}(d)}\log^{2}{n}/\log\log{n}$ bits per vertex for every $n$-point metric with doubling dimension $d$. (See Section˜5.)
{restatable}
theoremroutingUBdoubling For every $n$ and every $n$-point metric with doubling dimension $d$, there is a 3-hop routing scheme with stretch $(1+\epsilon)$ that uses $\epsilon^{-\tilde{O}(d)}\cdot\log^{2}{n}/\log\log{n}$ bits per vertex.
This provides the first routing scheme in Euclidean and doubling metrics, where the number of hops is $o(\log n)$, and the labels consist of $o(\log^{2}{n})$ bits.
Arboricity.
In Section˜4 we focus on low-arboricity shortcuttings. We show how to construct shortcuttings of hierarchically separated trees (HSTs) where the arboricity grows proportionally to an inverse Ackermann function. Our starting point is a sparse shortcutting of an $n$-vertex path. We use the well-known fact about arboricity — if the edges of a given graph can be oriented so that the in-degree is bounded by $\beta$, then the arboricity is at most $\beta+1$. Our key observation is that one can split every edge $(u,v)$ of a sparse shortcutting into two edges $(u,w)$ and $(w,v)$ in order to reduce the in-degrees of $u$ and $v$. By carefully choosing vertex $w$ for every edge $(u,v)$, we show that the arboricity can be bounded by $\alpha_{k/2+1}(n)$ when the hop-diameter is $k$. Roughly speaking, we achieve the same sparsity bound with a twice as large hop-diameter, which is the consequence of splitting each edge into two. Our construction for paths is then readily applicable to HSTs. Using the known HST cover [FL22a], we obtain the following theorem. See Section˜4 for details.
{restatable}
theoremarbUBdoubling Let $k$ be an even integer and let $\epsilon\in(0,1/6)$ be an arbitrary parameter. Then, for every positive integer $n$, every $n$-point metric with doubling dimension $d$ admits a $(1+\epsilon)$-spanner with hop-diameter $k$ and arboricity $\epsilon^{-O(d)}\alpha_{k/2+1}(n)$.
This significantly strengthens the construction of Arya et al. [ADM+95], by providing a uniformly sparse (rather than just sparse) construction with constant hop-diameter: For hop-diameter $k$, we transition from sparsity $\epsilon^{-O(d)}\alpha_{k}(n)$ to arboricity $\epsilon^{-O(d)}\alpha_{k/2+1}(n)$. In particular, we get arboricity $O(\log^{*}n)$ with a hop-diameter of 6. Recall that, in contrast to arboricity, one cannot achieve similar dependencies on the treewidth.
For general trees, we obtain a $k$-hop shortcutting with arboricity $O(\log^{12/(k+4)}{n})$. Note that this is an improvement over the treewidth bound given in Section˜1.1. Specifically, when $k=\Theta(\log\log n)$, the arboricity is $O(1)$, while the treewidth is $\Omega(\log\log n)$. Using this result and the known tree cover constructions [CCL+23, CCL+24a, BFN19, MN07], we obtain low-hop spanners for planar, minor-free, and general metrics with small arboricity. See Section˜4 for details.